# Inference-Time Alignment for Diffusion Models via Doob’s Matching

### v1

@@ D.2  Convergence rate of the controllable diffusion models
(D.17) $\leq\frac{2(4d\mu_{T_{0}}^{2}+8\sigma_{T_{0}}^{2})\log(\varepsilon^{-1})}{\mu_{T_{0}}^{2}}\frac{C^{\prime}\varepsilon^{3}}{\sigma_{T_{0}}^{4}},$
@@
where $C$ is a constant depending only on $d$, $\bar{B}$, and $\underline{B}$, and the second inequality holds from Corollary D.1 with $\delta=\varepsilon^{3}$.
Step 4. Conclusion. Substituting (D.12), (D.13), and (D.17) into (D.11) yields
@@
${\mathcal{W}}_{2}^{2}(q_{0},({\mathcal{M}}\circ{\mathcal{T}}_{R})_{\sharp}\widehat{q}_{T-T_{0}})$ $\lesssim\frac{d\sigma_{T_{0}}^{2}}{\mu_{T_{0}}^{2}}+\frac{d2^{d}}{\mu_{T_{0}}^{2}}\varepsilon+\frac{2(4d\mu_{T_{0}}^{2}+8\sigma_{T_{0}}^{2})\log(\varepsilon^{-1})}{\mu_{T_{0}}^{2}}\frac{C^{\prime}\varepsilon^{3}}{\sigma_{T_{0}}^{4}}$
$\leq C\Big\{\sigma_{T_{0}}^{2}+\varepsilon+\frac{\varepsilon^{3}}{\sigma_{T_{0}}^{4}}\log\Big(\frac{1}{\varepsilon}\Big)\Big\},$
@@
where $C$ is a constant depending only on $d$, $\bar{B}$, and $\underline{B}$. Letting $\sigma_{T_{0}}^{2}=\Theta(\varepsilon)$, i.e., $T_{0}=\Theta(\varepsilon)$, completes the proof. ∎
Appendix E Auxilary Lemmas
Lemma E.1.
@@ Appendix E Auxilary Lemmas
(E.1) ${\mathbb{E}}\Big[\exp\Big(\frac{\|\mu_{t}\mathbf{X}_{0}\|_{2}^{2}}{2d\mu_{t}^{2}}\Big)\Big]\leq 2.$
@@
(E.2) $=(2\pi)^{-\frac{d}{2}}\int\exp\Big(-\frac{\|{\bm{\varepsilon}}\|_{2}^{2}}{4}\Big)\,\mathrm{d}{\bm{\varepsilon}}\leq 2^{d}.$
@@
(E.3) $\leq{\mathbb{E}}\Big[\exp\Big(\frac{\|\mu_{t}\mathbf{X}_{0}\|_{2}^{2}}{2d\mu_{t}^{2}}\Big)\Big]{\mathbb{E}}\Big[\exp\Big(\frac{\|\sigma_{t}{\bm{\varepsilon}}\|_{2}^{2}}{4\sigma_{t}^{2}}\Big)\Big]\leq 2^{d+1},$
where the the first inequality follows from Cauchy-Schwarz inequality, the second inequality holds from the independence of $\mathbf{X}_{0}$ and $\bm{\varepsilon}$}, and the last inequality is due to (E.1) and (E.2). Then we aim to bound the tail probability. For each $\xi>0$, we have
@@
where the first inequality invokes Markov’s inequality, and the last inequality is due to (E.3). This completes the proof. ∎
Corollary E.2.
@@
Proof of Corollary E.2.
Under Assumptions 1 and 2, $\mathrm{supp}(q_{0})=\mathrm{supp}(p_{0})$. The same argument as Lemma E.1 completes the proof. ∎
Lemma E.3.

### v2

@@ E.2  Convergence rate of the controllable diffusion models
(E.16) $\leq\frac{2(4d\mu_{T_{0}}^{2}+8\sigma_{T_{0}}^{2})\log(\varepsilon^{-1})}{\mu_{T_{0}}^{2}}\frac{C^{\prime}\varepsilon^{3}}{\sigma_{T_{0}}^{4}}\log^{2}\Big(\frac{1}{\varepsilon}\Big),$
@@
where $C$ is a constant depending only on $d$, $\bar{B}$, and $\underline{B}$, and the second inequality holds from Corollary E.1 with $\delta=\varepsilon^{3}$.
Step 4. Conclusion. Substituting (E.11), (E.12), and (E.16) into (E.10) yields
@@
${\mathcal{W}}_{2}^{2}(q_{0},({\mathcal{M}}\circ{\mathcal{T}}_{R})_{\sharp}\widehat{q}_{T-T_{0}})$ $\lesssim\frac{d\sigma_{T_{0}}^{2}}{\mu_{T_{0}}^{2}}+\frac{d2^{d}}{\mu_{T_{0}}^{2}}\varepsilon+\frac{2(4d\mu_{T_{0}}^{2}+8\sigma_{T_{0}}^{2})\log(\varepsilon^{-1})}{\mu_{T_{0}}^{2}}\frac{C^{\prime}\varepsilon^{3}}{\sigma_{T_{0}}^{4}}\log^{2}\Big(\frac{1}{\varepsilon}\Big)$
$\leq C\Big\{\sigma_{T_{0}}^{2}+\varepsilon+\frac{\varepsilon^{3}}{\sigma_{T_{0}}^{4}}\log^{3}\Big(\frac{1}{\varepsilon}\Big)\Big\},$
@@
where $C$ is a constant depending only on $d$, $\bar{B}$, and $\underline{B}$. Letting $\sigma_{T_{0}}^{2}\asymp\varepsilon$, i.e., $T_{0}\asymp\varepsilon$, completes the proof. ∎
Appendix F Derivations in Section 5.4
Proposition 5.7.
Suppose Assumptions 4 and 2 hold. Then for any $t\in(0,T)$ and $\mathbf{x}\in{\mathbb{R}}^{d}$, we have
(F.1) $h_{t}^{*}(\mathbf{x})=\bar{h}_{t}^{*}(\mathbf{P}^{\top}\mathbf{x})\coloneq{\mathbb{E}}[w(\mathbf{P}\bar{\mathbf{X}}_{T}^{\leftarrow})\mid\bar{\mathbf{X}}_{t}^{\leftarrow}=\mathbf{P}^{\top}\mathbf{x}].$
Proof of Proposition 5.7.
According to Assumption 4, a particle $\mathbf{X}_{0}$ following $p_{0}$ satisfies
$\mathbf{X}_{0}\stackrel{{\scriptstyle\,\mathrm{d}}}{{=}}\mathbf{P}\bar{\mathbf{X}}_{0},\quad\bar{\mathbf{X}}_{0}\sim\bar{p}_{0}.$
We first establish the relations between $p_{t}$ and $\bar{p}_{t}$. It is straightforward that
$p_{t}(\mathbf{x})$ $=\int\varphi_{d}(\mathbf{x};\mu_{t}\mathbf{x}_{0},\sigma_{t}^{2}\mathbf{I}_{d})p_{0}(\mathbf{x}_{0})\,\mathrm{d}\mathbf{x}_{0}$
$=\int\varphi_{d}(\mathbf{x};\mu_{t}\mathbf{x}_{0},\sigma_{t}^{2}\mathbf{I}_{d})\Big(\int\delta_{\mathbf{P}\bar{\mathbf{x}}_{0}}(\mathbf{x}_{0})\bar{p}_{0}(\bar{\mathbf{x}}_{0})\,\mathrm{d}\bar{\mathbf{x}}_{0}\Big)\,\mathrm{d}\mathbf{x}_{0}$
$=\int\Big(\int\varphi_{d}(\mathbf{x};\mu_{t}\mathbf{x}_{0},\sigma_{t}^{2}\mathbf{I}_{d})\delta_{\mathbf{P}\bar{\mathbf{x}}_{0}}(\mathbf{x}_{0})\,\mathrm{d}\mathbf{x}_{0}\Big)\bar{p}_{0}(\bar{\mathbf{x}}_{0})\,\mathrm{d}\bar{\mathbf{x}}_{0}$
$=\int\varphi_{d}(\mathbf{x};\mu_{t}\mathbf{P}\bar{\mathbf{x}}_{0},\sigma_{t}^{2}\mathbf{I}_{d})\bar{p}_{0}(\bar{\mathbf{x}}_{0})\,\mathrm{d}\bar{\mathbf{x}}_{0}$
$=(2\pi\sigma_{t}^{2})^{-\frac{d}{2}}\int\exp\Big(-\frac{\|\mathbf{x}-\mu_{t}\mathbf{P}\bar{\mathbf{x}}_{0}\|_{2}^{2}}{2\sigma_{t}^{2}}\Big)\bar{p}_{0}(\bar{\mathbf{x}}_{0})\,\mathrm{d}\bar{\mathbf{x}}_{0}$
$=(2\pi\sigma_{t}^{2})^{-\frac{d}{2}}\int\exp\Big(-\frac{\|(\mathbf{I}_{d}-\mathbf{P}\mathbf{P}^{\top})\mathbf{x}+\mathbf{P}\mathbf{P}^{\top}\mathbf{x}-\mu_{t}\mathbf{P}\bar{\mathbf{x}}_{0}\|_{2}^{2}}{2\sigma_{t}^{2}}\Big)\bar{p}_{0}(\bar{\mathbf{x}}_{0})\,\mathrm{d}\bar{\mathbf{x}}_{0}$
$=(2\pi\sigma_{t}^{2})^{-\frac{d}{2}}\int\exp\Big(-\frac{\|(\mathbf{I}_{d}-\mathbf{P}\mathbf{P}^{\top})\mathbf{x}\|_{2}^{2}+\|\mathbf{P}\mathbf{P}^{\top}\mathbf{x}-\mu_{t}\mathbf{P}\bar{\mathbf{x}}_{0}\|_{2}^{2}}{2\sigma_{t}^{2}}\Big)\bar{p}_{0}(\bar{\mathbf{x}}_{0})\,\mathrm{d}\bar{\mathbf{x}}_{0}$
$=(2\pi\sigma_{t}^{2})^{-\frac{d}{2}}\exp\Big(-\frac{\|(\mathbf{I}_{d}-\mathbf{P}\mathbf{P}^{\top})\mathbf{x}\|_{2}^{2}}{2\sigma_{t}^{2}}\Big)\int\exp\Big(-\frac{\|\mathbf{P}^{\top}\mathbf{x}-\mu_{t}\bar{\mathbf{x}}_{0}\|_{2}^{2}}{2\sigma_{t}^{2}}\Big)\bar{p}_{0}(\bar{\mathbf{x}}_{0})\,\mathrm{d}\bar{\mathbf{x}}_{0}$
(F.2) $=\exp\Big(-\frac{\|(\mathbf{I}_{d}-\mathbf{P}\mathbf{P}^{\top})\mathbf{x}\|_{2}^{2}}{2\sigma_{t}^{2}}\Big)\bar{p}_{t}(\mathbf{P}^{\top}\mathbf{x}),$
where the seventh equality invokes the fact that $(\mathbf{I}_{d}-\mathbf{P}\mathbf{P}^{\top})\mathbf{x}$ is orthogonal to $\mathbf{P}\mathbf{P}^{\top}\mathbf{x}-\mu_{t}\mathbf{P}\bar{\mathbf{x}}_{0}$, the eighth equality is due to $\|\mathbf{P}\mathbf{v}\|_{2}=\|\mathbf{v}\|_{2}$ for each $\mathbf{v}\in{\mathbb{R}}^{d}$, and the last equality used (5.5). Then by a similar argument as the density, we find
$h_{T-t}^{*}(\mathbf{x})={\mathbb{E}}[w(\mathbf{X}_{0})\mid\mathbf{X}_{t}=\mathbf{x}]$
$=\frac{1}{p_{t}(\mathbf{x})}\int w(\mathbf{x}_{0})\varphi_{d}(\mathbf{x};\mu_{t}\mathbf{x}_{0},\sigma_{t}^{2}\mathbf{I}_{d})p_{0}(\mathbf{x}_{0})\,\mathrm{d}\mathbf{x}_{0}$
$=\frac{1}{p_{t}(\mathbf{x})}\int w(\mathbf{P}\bar{\mathbf{x}}_{0})\varphi_{d}(\mathbf{x};\mu_{t}\mathbf{P}\bar{\mathbf{x}}_{0},\sigma_{t}^{2}\mathbf{I}_{d})\bar{p}_{0}(\bar{\mathbf{x}}_{0})\,\mathrm{d}\bar{\mathbf{x}}_{0}$
$=\frac{1}{p_{t}(\mathbf{x})}(2\pi\sigma_{t}^{2})^{-\frac{d}{2}}\int w(\mathbf{P}\bar{\mathbf{x}}_{0})\exp\Big(-\frac{\|\mathbf{x}-\mu_{t}\mathbf{P}\bar{\mathbf{x}}_{0}\|_{2}^{2}}{2\sigma_{t}^{2}}\Big)\bar{p}_{0}(\bar{\mathbf{x}}_{0})\,\mathrm{d}\bar{\mathbf{x}}_{0}$
$=\frac{1}{p_{t}(\mathbf{x})}(2\pi\sigma_{t}^{2})^{-\frac{d}{2}}\int w(\mathbf{P}\bar{\mathbf{x}}_{0})\exp\Big(-\frac{\|(\mathbf{I}_{d}-\mathbf{P}\mathbf{P}^{\top})\mathbf{x}+\mathbf{P}\mathbf{P}^{\top}\mathbf{x}-\mu_{t}\mathbf{P}\bar{\mathbf{x}}_{0}\|_{2}^{2}}{2\sigma_{t}^{2}}\Big)\bar{p}_{0}(\bar{\mathbf{x}}_{0})\,\mathrm{d}\bar{\mathbf{x}}_{0}$
$=\frac{1}{p_{t}(\mathbf{x})}\exp\Big(-\frac{\|(\mathbf{I}_{d}-\mathbf{P}\mathbf{P}^{\top})\mathbf{x}\|_{2}^{2}}{2\sigma_{t}^{2}}\Big)\int w(\mathbf{P}\bar{\mathbf{x}}_{0})\varphi_{d}(\mathbf{P}^{\top}\mathbf{x};\mu_{t}\bar{\mathbf{x}}_{0},\sigma_{t}^{2}\mathbf{I}_{d})\bar{p}_{0}(\bar{\mathbf{x}}_{0})\,\mathrm{d}\bar{\mathbf{x}}_{0}$
$=\frac{1}{\bar{p}_{t}(\mathbf{P}^{\top}\mathbf{x})}\int w(\mathbf{P}\bar{\mathbf{x}}_{0})\varphi_{d}(\mathbf{P}^{\top}\mathbf{x};\mu_{t}\bar{\mathbf{x}}_{0},\sigma_{t}^{2}\mathbf{I}_{d})\bar{p}_{0}(\bar{\mathbf{x}}_{0})\,\mathrm{d}\bar{\mathbf{x}}_{0}$
$={\mathbb{E}}[w(\mathbf{P}\bar{\mathbf{X}}_{0})\mid\bar{\mathbf{X}}_{t}=\mathbf{P}^{\top}\mathbf{x}]={\mathbb{E}}[w(\mathbf{P}\bar{\mathbf{X}}_{T}^{\leftarrow})\mid\bar{\mathbf{X}}_{T-t}^{\leftarrow}=\mathbf{P}^{\top}\mathbf{x}],$
where the second and the eighth equalities are due to Bayes’ rule, the seventh equality holds from (F.2). This completes the proof. ∎
Proposition 5.8.
Suppose Assumptions 4 and 2 hold. Then for all $t\in(0,T)$ and $\bar{\mathbf{x}}\in{\mathbb{R}}^{d^{*}}$, the following bounds hold:
  1. (i)
    $\underline{B}\leq\bar{h}_{t}^{*}(\bar{\mathbf{x}})\leq\bar{B}$;
  2. (ii)
    $\max_{1\leq k\leq d}|D_{k}\bar{h}_{t}^{*}(\bar{\mathbf{x}})|\leq 2\sigma_{T-t}^{-2}\bar{B}$; and
  3. (iii)
    $\max_{1\leq k,\ell\leq d}|D_{k\ell}^{2}\bar{h}_{t}^{*}(\bar{\mathbf{x}})|\leq 6\sigma_{T-t}^{-4}\bar{B}$,
Proof of Proposition 5.8.
By the simialr argument as Lemmas C.1, C.4, and C.5, we conclude the desired results. ∎
By a similar argument as Lemma D.8, we have the following approximation error bounds for low-dimensional Doob’s $h$-function $\bar{h}_{t}^{*}$ (5.7).
Lemma F.1 (Approximation error).
Suppose Assumptions 4 and 2 hold. Let $R\geq 1$, and let the hypothesis class ${\mathscr{H}}_{t}$ be defined as (5.8) with $L\leq C\log N$ and $S\leq N^{d^{*}}$, then
$\|h_{t}-\bar{h}_{t}^{*}\|_{L^{2}(p_{T-t})}^{2}$ $\leq C\frac{\bar{B}^{2}\log^{4}N}{\sigma_{T-t}^{8}N^{4}},$
$\|\nabla h_{t}-\nabla\bar{h}_{t}^{*}\|_{L^{2}(p_{T-t})}^{2}$ $\leq C\frac{\bar{B}^{2}\log^{2}N}{\sigma_{T-t}^{8}N^{2}}.$
provided that $R^{2}=(4d\mu_{t}^{2}+8\sigma_{t}^{2})\log N^{4}$, where $C$ is a constant only depending on $d^{*}$.
Theorem 5.9.
(F.3) ${\mathscr{H}}_{t}\coloneq\left\{h_{t}\in N(L,S):\begin{aligned} &\sup_{\mathbf{x}\in{\mathbb{R}}^{d}}h_{t}(\mathbf{x})\leq\bar{B},\penalty 10000\ \inf_{\mathbf{x}\in{\mathbb{R}}^{d}}h_{t}(\mathbf{x})\geq\underline{B},\\ &\max_{1\leq k\leq d}\sup_{\mathbf{x}\in{\mathbb{R}}^{d}}|D_{k}h_{t}(\mathbf{x})|\leq 2\sigma_{T-t}^{-2}\bar{B}\end{aligned}\right\},$
where $L={\mathcal{O}}(\log n)$ and $S={\mathcal{O}}(n^{\frac{d^{*}}{d^{*}+8}})$. Let $\widehat{h}_{t}^{\lambda}$ be the gradient-regularized empirical risk minimizer defined as (4.6), and let $h_{t}^{*}$ be the Doob’s $h$-function defined as (3.10). Then the following inequality holds:
${\mathbb{E}}\Big[\|\nabla\log\widehat{h}_{t}^{\lambda}-\nabla\log h_{t}^{*}\|_{L^{2}(p_{T-t})}^{2}\Big]\leq C\sigma_{T-t}^{-8}n^{-\frac{2}{d^{*}+8}}\log^{4}n,$
provided that the regularization parameter $\lambda$ is set as $\lambda={\mathcal{O}}(n^{-\frac{2}{d^{*}+8}})$, where $C$ is a constant depending only on $d^{*}$, $\bar{B}$, and $\underline{B}$.
Proof of Theorem 5.9.
Using the same arguments as the proof of Theorem 5.3 and applying Lemma F.1 completes the proof. ∎
Appendix G Auxilary Lemmas
Lemma G.1.
@@ Appendix G Auxilary Lemmas
(G.1) ${\mathbb{E}}\Big[\exp\Big(\frac{\|\mu_{t}\mathbf{X}_{0}\|_{2}^{2}}{2d\mu_{t}^{2}}\Big)\Big]\leq 2.$
@@
(G.2) $=(2\pi)^{-\frac{d}{2}}\int\exp\Big(-\frac{\|{\bm{\varepsilon}}\|_{2}^{2}}{4}\Big)\,\mathrm{d}{\bm{\varepsilon}}\leq 2^{d}.$
@@
(G.3) $\leq{\mathbb{E}}\Big[\exp\Big(\frac{\|\mu_{t}\mathbf{X}_{0}\|_{2}^{2}}{2d\mu_{t}^{2}}\Big)\Big]{\mathbb{E}}\Big[\exp\Big(\frac{\|\sigma_{t}{\bm{\varepsilon}}\|_{2}^{2}}{4\sigma_{t}^{2}}\Big)\Big]\leq 2^{d+1},$
where the the first inequality follows from Cauchy-Schwarz inequality, the second inequality holds from the independence of $\mathbf{X}_{0}$ and $\bm{\varepsilon}$}, and the last inequality is due to (G.1) and (G.2). Then we aim to bound the tail probability. For each $\xi>0$, we have
@@
where the first inequality invokes Markov’s inequality, and the last inequality is due to (G.3). This completes the proof. ∎
Corollary G.2.
@@
Proof of Corollary G.2.
Under Assumptions 1 and 2, $\mathrm{supp}(q_{0})=\mathrm{supp}(p_{0})$. The same argument as Lemma G.1 completes the proof. ∎
Lemma G.3.